OpenStudy (1645323):

Solve by completing the square. x2- 2x - 4 = 0 Select one: a. -12 b. 15 c. -15 d. 12

OpenStudy (cwrw238):

first rearrange: x^2 - 2x = 4 (x - 1)^2 -1 = 4 (x-1)^2 = 5 can you continue do you understand how I got the second line?

OpenStudy (1645323):

i Got Square root 5+1

OpenStudy (cwrw238):

thats correct - well sqrt5 +1 , -sqrt5 + 1 the options are incorrect